# Tight frames over the quaternions and equiangular lines

### v2

@@ 3 From $\mathbb{R}$ to $\mathbb{C}$ and $\mathbb{C}$ to $\mathbb{H}$, and back
$[v]_{\mathbb{C}}=\begin{pmatrix}z\cr\overline{w}\end{pmatrix},\quad[vj]_{\mathbb{C}}=\begin{pmatrix}-w\cr\overline{z}\end{pmatrix},\qquad\langle[v]_{\mathbb{C}},[vj]_{\mathbb{C}}\rangle_{\mathbb{C}}=0.$
Thus $[\cdot]_{\mathbb{C}}$ maps $k$-dimensional $\mathbb{H}$-subspaces of $\mathbb{H}^{d}$ to $(2k)$-dimensional $\mathbb{C}$-subspaces of $\mathbb{C}^{2d}$.
Let $L:\mathbb{H}^{n}\to\mathbb{H}^{m}$ be an $\mathbb{H}$-linear map be represented as a $\mathbb{C}$-linear map $[L]_{\mathbb{C}}:\mathbb{C}^{2n}\to\mathbb{C}^{2m}$ under this identification, i.e., $[L]_{\mathbb{C}}:=[\cdot]_{\mathbb{C}}L[\cdot]_{\mathbb{C}}^{-1}$. In view of (3.28), its standard matrix $[L]_{\mathbb{H}}\in\mathbb{H}^{m\times n}$ has a unique decomposition
$[L]_{\mathbb{H}}=A+Bj,\qquad A,B\in\mathbb{C}^{m\times n}.$
@@
Let $[\cdot]_{\mathbb{C}}:\mathbb{H}^{d}\to\mathbb{C}^{2d}$ be the correspondence (3.29) between $\mathbb{H}^{d}$ and $\mathbb{C}^{2d}$. Then
@@
Proof: The sequence $V=V_{1}+V_{2}j$ is tight frame for $\mathbb{H}^{d}$ if and only if
@@
which is clearly satisfied if $V$ corresponds to a tight frame for $\mathbb{C}^{2d}$ (by Proposition 3.1).
@@
so that this in turn gives a tight frame for $\mathbb{R}^{4}$, i.e., $[e_{1},e_{3},e_{2},e_{4}]$.
@@
Consider the Gramian of the SIC of four vectors in $\mathbb{C}^{2}$ (Example 2.1). The contribution to $\|V^{*}V\|_{F}$ of the diagonal entries, which are all real, is $4$, and for the off diagonal entries it is $412{1\over 3}=$ 4. Thus the SIC corresponds to a tight frame for $\mathbb{R}^{4}$ if and only if its vectors can be scaled so that the off diagonal entries of the Gramian are pure imaginary. This can in fact be done, e.g., take $V=[v,iSv,i\Omega v,-S\Omega v]$, to obtain
@@
We now summarise some basic results about $[\cdot]_{\mathbb{F}}$, $\mathbb{F}=\mathbb{R},\mathbb{C}$, and the associated linear maps, in a unified form. We first observe that in the literature, there is some variation in the definitions, in particular, the ordering of $[v]_{\mathbb{R}}$ can be either of
$[v]_{\mathbb{R}}=\begin{pmatrix}\mathop{\rm Re}\nolimits(v)\cr\mathop{\rm Im}\nolimits(v)\end{pmatrix},\qquad\begin{pmatrix}\mathop{\rm Re}\nolimits(v_{1})\cr\mathop{\rm Im}\nolimits(v_{1})\cr\vdots\cr\mathop{\rm Re}\nolimits(v_{d})\cr\mathop{\rm Im}\nolimits(v_{d})\end{pmatrix},$
and similarly for $[v]_{\mathbb{C}}$. In the latter case (cf [Hog77], [Rod14] for $[v]_{\mathbb{C}}$), the matrix representation $[A]_{\mathbb{R}}$ is then obtained by replacing the entry $a_{jk}$ of the matrix $A$ by the matrix
@@
$[L]_{\mathbb{F}}=\begin{pmatrix}A&-B\cr\overline{B}&\overline{A}\end{pmatrix},\qquad[L^{*}]_{\mathbb{F}}=\begin{pmatrix}A^{*}&B^{T}\cr-B^{*}&A^{T}\end{pmatrix}=[L]_{\mathbb{F}}^{*},\qquad\mathbb{F}=\mathbb{R},\mathbb{C}.$ (3.36)
@@
The $\mathbb{F}$-linear maps $[\cdot]_{\mathbb{F}}$, $\mathbb{F}=\mathbb{R},\mathbb{C}$ given by (3.19) and (3.29) have the following properties
@@
    $[AB]_{\mathbb{F}}=[A]_{\mathbb{F}}[B]_{\mathbb{F}}$, $[\lambda A]_{\mathbb{F}}=\lambda[A]_{\mathbb{F}}$, $\lambda\in\mathbb{R}$, and $[A^{*}]_{\mathbb{F}}=[A]_{\mathbb{F}}^{*}$.
@@
    They map invertible linear maps to invertible linear maps, with $[A^{-1}]_{\mathbb{F}}=[A]_{\mathbb{F}}^{-1}$.
@@
$[\mathbb{H}^{m\times n}]_{\mathbb{C}}=\{A\in\mathbb{C}^{2m\times 2n}:J_{m}A=\overline{A}J_{n}\},\qquad J_{\ell}:=[jI_{\ell}]_{\mathbb{C}}=\begin{pmatrix}0&-I_{\ell}\cr I_{\ell}&0\end{pmatrix}.$
@@
If $G$ is a group of $d\times d$ matrices over $\mathbb{C}$ or $\mathbb{H}$, then it follows from Theorem 3.3 that $[G]_{\mathbb{F}}=\{[g]_{\mathbb{F}}:g\in G\}$ is an isomorphic group of $(2d)\times(2d)$ matrices. As an example, the quaternions $Q=\{\pm 1,\pm i,\pm j,\pm k\}$ are generated by $i$ and $j$, and so the groups of unitary matrices $[Q]_{\mathbb{C}}$ and $[[Q]_{\mathbb{C}}]_{\mathbb{R}}$ are generated by
$[i]_{\mathbb{C}}=\begin{pmatrix}i&0\cr 0&-i\end{pmatrix},\qquad[j]_{\mathbb{C}}=\begin{pmatrix}0&-1\cr 1&0\end{pmatrix},$
$[[i]_{\mathbb{C}}]_{\mathbb{R}}=\begin{pmatrix}0&0&-1&0\cr 0&0&0&1\cr 1&0&0&0\cr 0&-1&0&0\end{pmatrix},\qquad[[j]_{\mathbb{C}}]_{\mathbb{R}}=\begin{pmatrix}0&-1&0&0\cr 1&0&0&0\cr 0&0&0&-1\cr 0&0&1&0\end{pmatrix},\qquad$
respectively. These representations of $Q$ are well known.
@@
$[V]_{\mathbb{C}}=\bigl{[}[v_{1}]_{\mathbb{C}},\ldots,[v_{n}]_{\mathbb{C}},[v_{1}j]_{\mathbb{C}},\ldots,[v_{n}j]_{\mathbb{C}}\bigr{]}$
@@
$|\langle v_{a},v_{b}\rangle|^{2}=\lambda\hskip 10.00002pt\Longleftrightarrow\hskip 10.00002pt(V_{a}^{*}V_{b})^{*}(V_{a}^{*}V_{b})=\lambda I.$
Thus $(v_{a})$ gives a set of equiangular lines in $\mathbb{H}^{d}$ if and only if the the off diagonal entries of the block Gramian $[V_{1},\ldots,V_{n}]^{*}[V_{1},\ldots,V_{n}]$ are unitary matrices, up to a fixed scalar. An $n\times n$ block matrix with this structural form ($2\times 2$ blocks, positive semi-definite of rank $2d$), which corresponds to equi-isoclinic planes in $\mathbb{C}^{2d}$, can then be mapped back (under $[\cdot]_{\mathbb{C}}^{-1}$) to the Gramian of $n$ equiangular lines in $\mathbb{H}^{d}$, see Theorem 13, [ET20].
@@
$=\langle[v_{a}]_{\mathbb{C}},[v_{b}]_{\mathbb{C}}\rangle_{\mathbb{C}}\langle[v_{b}]_{\mathbb{C}},[v_{a}]_{\mathbb{C}}\rangle_{\mathbb{C}}+\langle[v_{a}j]_{\mathbb{C}},[v_{b}]_{\mathbb{C}}\rangle_{\mathbb{C}}\langle[v_{b}]_{\mathbb{C}},[v_{a}j]_{\mathbb{C}}\rangle_{\mathbb{C}}$
@@
$=\langle[v_{a}]_{\mathbb{C}},[v_{b}]_{\mathbb{C}}\rangle_{\mathbb{C}}\langle[v_{b}j]_{\mathbb{C}},[v_{a}]_{\mathbb{C}}\rangle_{\mathbb{C}}+\langle[v_{a}j]_{\mathbb{C}},[v_{b}]_{\mathbb{C}}\rangle_{\mathbb{C}}\langle[v_{b}j]_{\mathbb{C}},[v_{a}j]_{\mathbb{C}}\rangle_{\mathbb{C}}$
$=\mathop{\rm Co}\nolimits_{1}(\langle v_{a},v_{b}\rangle_{\mathbb{H}})(-\mathop{\rm Co}\nolimits_{2}(\langle v_{b},v_{a}\rangle_{\mathbb{H}}))-\mathop{\rm Co}\nolimits_{2}(\langle v_{a},v_{b}\rangle_{\mathbb{H}})\mathop{\rm Co}\nolimits_{1}(\langle v_{b}j,v_{a}j\rangle_{\mathbb{H}})$
$=\mathop{\rm Co}\nolimits_{1}(\langle v_{a},v_{b}\rangle_{\mathbb{H}})\mathop{\rm Co}\nolimits_{2}(\langle v_{a},v_{b}\rangle_{\mathbb{H}})-\mathop{\rm Co}\nolimits_{2}(\langle v_{a},v_{b}\rangle_{\mathbb{H}})\mathop{\rm Co}\nolimits_{1}(\langle v_{a},v_{b}\rangle_{\mathbb{H}})=0,$
@@
We consider the construction of $64$ equiangular lines in $\mathbb{C}^{8}$ by [Hog98]. These were obtained by finding $64$ unit vectors in $\mathbb{H}^{4}$ with angles ${1\over 9},{1\over 3}$ (as vertices of a quaternionic polytope). These were then mapped by $[\cdot]_{\mathbb{C}}$ to $64$ equiangular vectors in $\mathbb{C}^{8}$. We note that for $v,w\in\mathbb{H}^{d}$, $\alpha\in\mathbb{H}$, (3.30) gives
@@ 4 Group frames and $G$-matrices
Many tight frames of interest are the orbit one or more vectors under the unitary action of a group, e.g., the Weyl-Heisenberg SICs. There is a well developed theory of such group frames based in the theory of group representations (over $\mathbb{R}$ and $\mathbb{C}$) [VW05], [Wal13], [VW16], [Wal18]. We now give an indication of how this theory extends to representations over $\mathbb{H}$ (see [SS95]).
[…]
$S(hv)=\sum_{g\in G}gv\langle hv,gv\rangle=h\sum_{g\in G}h^{-1}gv\langle v,h^{-1}gv\rangle=hS(v),\quad g,h\in G,\ v\in\mathbb{H}^{d}.$ (4.38)
@@
The Gramian of a $G$-frame is a $G$-matrix, and conversely if the Gramian of a frame $(v_{g})_{g\in G}$ with vectors indexed by $G$ is a $G$-matrix, then it is a $G$-frame (adapt the proof of [Wal18] Theorem 10.3). An action (representation) of $G$ on $\mathbb{H}^{d}$ is irreducible if the only $G$-invariant subspaces of $\mathbb{H}^{d}$ are 00 and $\mathbb{H}^{d}$, i.e., $\mathop{\rm span}\nolimits_{\mathbb{H}}\{gv\}_{g\in G}=\mathbb{H}^{d}$, for all $v\neq 0$.
@@
$x={d\over|G|}{1\over\|v\|^{2}}\sum_{g\in G}gv\langle x,gv\rangle,\qquad\forall x\in\mathbb{H}^{d}.$
Proof: Fix $v\neq 0$, and let $S$ be the frame operator of $(gv)_{g\in G}$. Since $S$ is nonzero and positive semidefinite, it has an eigenvalue $\lambda>0$, with corresponding eigenvector $w$. By (4.38), $S$ commutes with the action of $g\in G$, we have
@@
has an irreducible unitary action on $\mathbb{H}^{2}$. It consists of all $128$ invertible matrices with two zero entries and two entries in $Q$. It contains the scalar matrices from $Q$ and its center is ±$\pm I$. Thus each orbit can be viewed as $16$ lines in $\mathbb{H}^{2}$ (as a left vector space). This is an example of (quaternionic) reflection group, i.e., a finite group generated by reflections (linear maps which act as the identity on a hyperplane). The finite irreducible quaternionic reflection groups have been classified (up to conjugacy) by Cohen [Coh80].
@@
(Quaternionic harmonic frames). The irreducible representations over $\mathbb{C}$ for an abelian group $G$ are all one-dimensional (this characterises abelian groups), and these “rows” of the character table are orthogonal, so by taking a set of rows of the character table one obtains a tight $G$-frame. Consider the quaternion group $G=Q$. This has four $1$-dimensional and one $2$-dimensional irreducible representations over $\mathbb{C}$. The $2$-dimensional absolutely irreducible representation splits into four $1$-dimensional representations over $\mathbb{H}$, corresponding to the outer automorphisms of the quaternions. In this way, one obtains a character table
$\begin{array}[]{lcccccccc}q\in Q&1&-1&i&-i&j&-j&k&-k\\ \hline\cr\chi_{1}&1&1&1&1&1&1&1&1\\ \chi_{2}&1&1&1&1&-1&-1&-1&-1\\ \chi_{3}&1&1&-1&-1&1&1&-1&-1\\ \chi_{4}&1&1&-1&-1&-1&-1&1&1\\ \chi_{5}&1&-1&i&-i&j&-j&k&-k\\ \chi_{6}&1&-1&j&-j&i&-i&-k&k\\ \chi_{7}&1&-1&-i&i&k&-k&j&-j\\ \chi_{8}&1&-1&k&-k&-i&i&-j&j\end{array}$
where the rows are orthogonal (cf [SS95]). Taking rows gives a $G$-frame. The columns of the character table are also orthogonal, so taking columns also gives a tight frame, but these are not $G$-frames, in general (as follows for abelian groups by Pontryagin duality). As an example, the frame obtained by taking the characters $\chi_{1}$ and $\chi_{5}$ (rows $1$ and $5$ of the character table) gives a (unit-norm) tight $Q$-frame for $\mathbb{H}^{2}$, with the inner products $\{1\pm 1,1\pm i,1\pm j,1\pm k\}$ occurring exactly once in every row (column) of the Gramian. This frame has two angles: each vector is orthogonal to one other, and makes a fixed angle with all the others. By comparison, taking columns $1$ and $3$ gives a unit-norm tight frame for $\mathbb{H}^{2}$, which is not a $G$-frame for any $G$, since its Gramian is not a $G$-matrix.

### v3

@@ 3 From $\mathbb{R}$ to $\mathbb{C}$ and $\mathbb{C}$ to $\mathbb{H}$, and back
$v]_{\mathbb{C}}=\begin{pmatrix}z\cr\overline{w}\end{pmatrix},\quad[vj]_{\mathbb{C}}=\begin{pmatrix}-w\cr\overline{z}\end{pmatrix},\qquad\langle[v]_{\mathbb{C}},[vj]_{\mathbb{C}}\rangle_{\mathbb{C}}=0.$
Thus $\cdot$$_{\mathbb{C}}$ maps $k$-dimensional $\mathbb{H}$-subspaces of $\mathbb{H}^{d}$ to $(2k)$-dimensional $\mathbb{C}$-subspaces of $\mathbb{C}^{2d}$.
Let $L:\mathbb{H}^{n}\to\mathbb{H}^{m}$ be an $\mathbb{H}$-linear map be represented as a $\mathbb{C}$-linear map $L]_{\mathbb{C}}:\mathbb{C}^{2n}\to\mathbb{C}^{2m}$ under this identification, i.e., $L]_{\mathbb{C}}:=[\cdot]_{\mathbb{C}}L[\cdot]_{\mathbb{C}}^{-1}$. In view of (3.29), its standard matrix $L]_{\mathbb{H}}\in\mathbb{H}^{m\times n}$ has a unique decomposition
$L]_{\mathbb{H}}=A+Bj,\qquad A,B\in\mathbb{C}^{m\times n}.$
@@
Let $\cdot]_{\mathbb{C}}:\mathbb{H}^{d}\to\mathbb{C}^{2d}$ be the correspondence (3.30) between $\mathbb{H}^{d}$ and $\mathbb{C}^{2d}$. Then
@@
Proof: The sequence $V=V_{1}+V_{2}j$ is a tight frame for $\mathbb{H}^{d}$ if and only if
@@
which is clearly satisfied if $V$ corresponds to a tight frame for $\mathbb{C}^{2d}$ (by Lemma 3.1).
@@
so that this in turn gives a tight frame for $\mathbb{R}^{4}$, i.e., $e_{1},e_{3},e_{2},e_{4}]$.
@@
Consider the Gramian of the SIC of four vectors in $\mathbb{C}^{2}$ (Example 2.1). The contribution to $\|V^{*}V\|_{F}$ of the diagonal entries, which are all real, is $4$, and for the off diagonal entries it is $12{1\over 3}=4$. Thus the SIC corresponds to a tight frame for $\mathbb{R}^{4}$ if and only if its vectors can be scaled so that the off diagonal entries of the Gramian are pure imaginary. This can in fact be done, e.g., take $V=[v,iSv,i\Omega v,-S\Omega v]$, to obtain
@@
We now summarise some basic results about $\cdot]_{\mathbb{F}}$, $\mathbb{F}=\mathbb{R},\mathbb{C}$, and the associated linear maps, in a unified form. We first observe that in the literature, there is some variation in the definitions, in particular, the ordering of $v]_{\mathbb{R}}$ can be either of
$v]_{\mathbb{R}}=\begin{pmatrix}\mathop{\rm Re}\nolimits(v)\cr\mathop{\rm Im}\nolimits(v)\end{pmatrix},\qquad\begin{pmatrix}\mathop{\rm Re}\nolimits(v_{1})\cr\mathop{\rm Im}\nolimits(v_{1})\cr\vdots\cr\mathop{\rm Re}\nolimits(v_{d})\cr\mathop{\rm Im}\nolimits(v_{d})\end{pmatrix}$,
and similarly for $v]_{\mathbb{C}}$. In the latter case (cf [Hog77], [Rod14] for $v]_{\mathbb{C}}$), the matrix representation $A]_{\mathbb{R}}$ is then obtained by replacing the entry $a_{jk}$ of the matrix $A$ by the matrix
@@
$L]_{\mathbb{F}}=\begin{pmatrix}A&-B\cr\overline{B}&\overline{A}\end{pmatrix},\qquad[L^{*}]_{\mathbb{F}}=\begin{pmatrix}A^{*}&B^{T}\cr-B^{*}&A^{T}\end{pmatrix}=[L]_{\mathbb{F}}^{*},\qquad\mathbb{F}=\mathbb{R},\mathbb{C}.$ (3.36)
@@
The $\mathbb{F}$-linear maps $\cdot]_{\mathbb{F}}$, $\mathbb{F}=\mathbb{R},\mathbb{C}$ given by (3.20) and (3.30) have the following properties
@@
    $AB]_{\mathbb{F}}=[A]_{\mathbb{F}}[B]_{\mathbb{F}}$, $\lambda A]_{\mathbb{F}}=\lambda[A]_{\mathbb{F}}$, $\lambda\in\mathbb{R}$, and $A^{*}]_{\mathbb{F}}=[A]_{\mathbb{F}}^{*}$.
@@
    They map invertible linear maps to invertible linear maps, with $A^{-1}]_{\mathbb{F}}=[A]_{\mathbb{F}}^{-1}$.
@@
$\mathbb{H}^{m\times n}]_{\mathbb{C}}=\{A\in\mathbb{C}^{2m\times 2n}:J_{m}A=\overline{A}J_{n}\},\qquad J_{\ell}:=[jI_{\ell}]_{\mathbb{C}}=\begin{pmatrix}0&-I_{\ell}\cr I_{\ell}&0\end{pmatrix}$.
@@
If $G$ is a group of $d\times d$ matrices over $\mathbb{C}$ or $\mathbb{H}$, then it follows from Theorem 3.3 that $G]_{\mathbb{F}}=\{[g]_{\mathbb{F}}:g\in G\}$ is an isomorphic group of $(2d)\times(2d)$ matrices. As an example, the quaternions $Q_{8}=\{\pm 1,\pm i,\pm j,\pm k\}$ are generated by $i$ and $j$, and so the groups of unitary matrices $Q_{8}]_{\mathbb{C}}$ and $Q_{8}]_{\mathbb{C}}]_{\mathbb{R}}$ are generated by
$i]_{\mathbb{C}}=\begin{pmatrix}i&0\cr 0&-i\end{pmatrix},\qquad[j]_{\mathbb{C}}=\begin{pmatrix}0&-1\cr 1&0\end{pmatrix}$,
$i]_{\mathbb{C}}]_{\mathbb{R}}=\begin{pmatrix}0&0&-1&0\cr 0&0&0&1\cr 1&0&0&0\cr 0&-1&0&0\end{pmatrix},\qquad[[j]_{\mathbb{C}}]_{\mathbb{R}}=\begin{pmatrix}0&-1&0&0\cr 1&0&0&0\cr 0&0&0&-1\cr 0&0&1&0\end{pmatrix}$$\qquad$
respectively. These representations of $Q_{8}$ are well known.
@@
$V]_{\mathbb{C}}=\bigl{[}[v_{1}]_{\mathbb{C}},\ldots,[v_{n}]_{\mathbb{C}},[v_{1}j]_{\mathbb{C}},\ldots,[v_{n}j]_{\mathbb{C}}\bigr{]}$
@@
$|\langle v_{a},v_{b}\rangle|^{2}=\lambda\quad\Longleftrightarrow\quad(V_{a}^{*}V_{b})^{*}(V_{a}^{*}V_{b})=\lambda I.$
Thus $(v_{a})$ gives a set of equiangular lines in $\mathbb{H}^{d}$ if and only if the off diagonal entries of the block Gramian $V_{1},\ldots,V_{n}]^{*}[V_{1},\ldots,V_{n}]$ are unitary matrices, up to a fixed scalar. An $n\times n$ block matrix with this structural form ($2\times 2$ blocks, positive semi-definite of rank $2d$), which corresponds to equi-isoclinic planes in $\mathbb{C}^{2d}$, can then be mapped back (under $\cdot]_{\mathbb{C}}^{-1}$) to the Gramian of $n$ equiangular lines in $\mathbb{H}^{d}$, see Theorem 13, [ET20].
@@
$=\langle[v_{b}]_{\mathbb{C}},[v_{a}]_{\mathbb{C}}\rangle_{\mathbb{C}}\langle[v_{a}]_{\mathbb{C}},[v_{b}]_{\mathbb{C}}\rangle_{\mathbb{C}}+\langle[v_{b}]_{\mathbb{C}},[v_{a}j]_{\mathbb{C}}\rangle_{\mathbb{C}}\langle[v_{a}j]_{\mathbb{C}},[v_{b}]_{\mathbb{C}}\rangle_{\mathbb{C}}$
@@
$=\langle[v_{b}]_{\mathbb{C}},[v_{a}]_{\mathbb{C}}\rangle_{\mathbb{C}}\langle[v_{a}]_{\mathbb{C}},[v_{b}j]_{\mathbb{C}}\rangle_{\mathbb{C}}+\langle[v_{b}]_{\mathbb{C}},[v_{a}j]_{\mathbb{C}}\rangle_{\mathbb{C}}\langle[v_{a}j]_{\mathbb{C}},[v_{b}j]_{\mathbb{C}}\rangle_{\mathbb{C}}$
$=\mathop{\rm Co}\nolimits_{1}(\langle v_{b},v_{a}\rangle_{\mathbb{H}})(-\mathop{\rm Co}\nolimits_{2}(\langle v_{a},v_{b}\rangle_{\mathbb{H}}))-\mathop{\rm Co}\nolimits_{2}(\langle v_{b},v_{a}\rangle_{\mathbb{H}})\mathop{\rm Co}\nolimits_{1}(\langle v_{a}j,v_{b}j\rangle_{\mathbb{H}})$
$=-\mathop{\rm Co}\nolimits_{1}(\langle v_{b},v_{a}\rangle_{\mathbb{H}})\mathop{\rm Co}\nolimits_{2}(\langle v_{a},v_{b}\rangle_{\mathbb{H}})+\mathop{\rm Co}\nolimits_{2}(\langle v_{b},v_{a}\rangle_{\mathbb{H}})\mathop{\rm Co}\nolimits_{1}(\langle v_{b},v_{a}\rangle_{\mathbb{H}})=0,$
@@
We consider the construction of $64$ equiangular lines in $\mathbb{C}^{8}$ by [Hog98]. These were obtained by finding $64$ unit vectors in $\mathbb{H}^{4}$ with angles ${1\over 9},{1\over 3}$ (as vertices of a quaternionic polytope). These were then mapped by $\cdot$$_{\mathbb{C}}$ to $64$ equiangular vectors in $\mathbb{C}^{8}$. We note that for $v,w\in\mathbb{H}^{d}$, $\alpha\in\mathbb{H}$, (3.31) gives
@@ 4 Group frames and $G$-matrices
Many tight frames of interest are the orbit of one or more vectors under the unitary action of a group, e.g., the Weyl-Heisenberg SICs. There is a well developed theory of such group frames based in the theory of group representations (over $\mathbb{R}$ and $\mathbb{C}$) [VW05], [Wal13], [VW16], [Wal18]. We now give an indication of how this theory extends to representations over $\mathbb{H}$ (see [SS95]).
A representation of a finite abstract group $G$ on $\mathbb{H}^{d}$ is a group homomorphism $\rho:G\to\mathop{\it GL}\nolimits(\mathbb{H}^{d})$ from $G$ to the invertible $d\times d$ matrices over $\mathbb{H}$, with equivalence defined in the usual way. We will consider only unitary representations, i.e.. those where the matrices $\rho(g)$ are unitary. For these, we will write the unitary action as $gv:=\rho(g)v$, and we note that $g^{*}v=g^{-1}v$. A frame (sequence of vectors) of the form $(gv)_{g\in G}$ is said to be a group frame (or $G$-frame) [Wal20a]. The frame operator of a group frame $(gv)_{g\in G}$ commutes with the frame operator, i.e.,
$S(hv)=\sum_{g\in G}gv\langle gv,hv\rangle=h\sum_{g\in G}h^{-1}gv\langle h^{-1}gv,v\rangle=hS(v),\quad h\in G,\ v\in\mathbb{H}^{d}.$ (4.38)
@@
The Gramian of a $G$-frame is a $G$-matrix, and conversely if the Gramian of a frame $(v_{g})_{g\in G}$ with vectors indexed by $G$ is a $G$-matrix, then it is a $G$-frame (adapt the proof of [Wal18] Theorem 10.3). An action (representation) of $G$ on $\mathbb{H}^{d}$ is irreducible if the only $G$-invariant subspaces of $\mathbb{H}^{d}$ are 0 and $\mathbb{H}^{d}$, i.e., $\mathop{\rm span}\nolimits_{\mathbb{H}}\{gv\}_{g\in G}=\mathbb{H}^{d}$, for all $v\neq 0$.
@@
$x={d\over|G|}{1\over\|v\|^{2}}\sum_{g\in G}gv\langle gv,x\rangle,\qquad\forall x\in\mathbb{H}^{d}.$
Proof: Fix $v\neq 0$, and let $S$ be the frame operator of $(gv)_{g\in G}$. Since $S$ is nonzero and positive semidefinite, it has an eigenvalue $\lambda>0$, with corresponding eigenvector $w$. By (4.38), $S$ commutes with the action of $g\in G$, so that
@@
has an irreducible unitary action on $\mathbb{H}^{2}$. It consists of all $128$ invertible matrices with two zero entries and two entries in $Q_{8}$. It contains the scalar matrices from $Q_{8}$ and its center is $\pm I$. Thus each orbit can be viewed as $16$ lines in $\mathbb{H}^{2}$ (as a left vector space). This is an example of a (quaternionic) reflection group, i.e., a finite group generated by reflections (linear maps which act as the identity on a hyperplane). The finite irreducible quaternionic reflection groups have been classified (up to conjugacy) by Cohen [Coh80].
@@
(Quaternionic harmonic frames). The irreducible representations over $\mathbb{C}$ for an abelian group $G$ are all one-dimensional (this characterises abelian groups), and these “rows” of the character table are orthogonal, so by taking a set of rows of the character table one obtains a tight $G$-frame. Consider the quaternion group $G=Q_{8}$. This has four $1$-dimensional and one $2$-dimensional irreducible representations over $\mathbb{C}$. The $2$-dimensional absolutely irreducible representation splits into four $1$-dimensional representations over $\mathbb{H}$, corresponding to the outer automorphisms of the quaternions. In this way, one obtains a character table
$\begin{array}[]{lcccccccc}q\in Q_{8}&1&-1&i&-i&j&-j&k&-k\\ \hline\cr\chi_{1}&1&1&1&1&1&1&1&1\\ \chi_{2}&1&1&1&1&-1&-1&-1&-1\\ \chi_{3}&1&1&-1&-1&1&1&-1&-1\\ \chi_{4}&1&1&-1&-1&-1&-1&1&1\\ \chi_{5}&1&-1&i&-i&j&-j&k&-k\\ \chi_{6}&1&-1&j&-j&i&-i&-k&k\\ \chi_{7}&1&-1&-i&i&k&-k&j&-j\\ \chi_{8}&1&-1&k&-k&-i&i&-j&j\end{array}$
where the rows are orthogonal (cf [SS95]). Taking rows gives a $G$-frame. The columns of the character table are also orthogonal, so taking columns also gives a tight frame, but these are not $G$-frames, in general (as follows for abelian groups by Pontryagin duality). As an example, the frame obtained by taking the characters $\chi_{1}$ and $\chi_{5}$ (rows $1$ and $5$ of the character table) gives a (unit-norm) tight $Q_{8}$-frame for $\mathbb{H}^{2}$, with the inner products $\{1\pm 1,1\pm i,1\pm j,1\pm k\}$ occurring exactly once in every row (column) of the Gramian. This frame has two angles: each vector is orthogonal to one other, and makes a fixed angle with all the others.